Ben can do a job in 12hours. Jake can do the same job in 16hours. Ben works for 9hours and leave. How many hours will it take jake to complete remaining work?
Answer by lwsshak3(11628)

Ben can do a job in 12hours. Jake can do the same job in 16hours. Ben works for 9hours and leave. How many hours will it take jake to complete remaining work?
***
1/12 =Ben"s work rate
1-16=Jake's work rate
let x=hours Jake must work to finish job
9(1/12)=x(1/16)=1
x/16=1-9/12=1/4
x=16/4=4
How many hours will it take jake to complete remaining work? 4